# Antiferromagnetism and spin excitations in a two-dimensional non-Hermitian Hatano-Nelson flux model

### v1

@@ I Introduction
In recent years one of the most popular avenues to study the evolution of open quantum-mechanical systems became its description via effective non-Hermitian Hamiltonians, which break unitarity [1]. At the origin of these research activities lies the seminal work on parity-time ($\mathcal{PT}$) symmetric Hamiltonian with balanced dissipation (loss) and gain energy terms added to Hermitian parent Hamiltonians [2, 3]. $\mathcal{PT}$ symmetry allows to construct further non-Hermitian Hamiltonians, which for some parameter space yield real eigenvalues and the transition between purely real and complex-valued (dissipative) energy spectrum is separated by exceptional points [4, 5]. The field of non-Hermitian systems has grown in recent years, driven by theoretical predictions and experimental discoveries across various field of physics such as optics, acoustics, ultra-cold atomic gases, and superconducting qubits [1, 6, 7]. In particular, non-Hermitian Hamiltonians have recently become quite promising due to the possibility of different realizations, for example using ultracold atoms [8, 9, 10]. The characteristic physical phenomena of non-Hermitian systems, such as formation of exceptional points, chiral transport around them, and the non-Hermitian skin effect are typically understood within a single-particle Hamiltonian formalism, and currently the interest shifts towards many-body systems where the interplay between interactions and non-Hermiticity plays a central role.
One prominent example of non-Hermiticity appears in the celebrated Hatano-Nelson (HN) model [11, 12, 13], which consists of a one-dimensional (1D) lattice with asymmetric (non-reciprocal) hoppings, responsible for the localization of all the bulk states at the edges, which leads to the emergence of the so-called topological non-Hermitian skin effect. The non-reciprocity of the hopping may be considered as resulting from some transverse pseudomagnetic field in a cylinder geometry [11]. By now, there has been studies of this model in the presence of interactions [14], disorder [15, 16], as well as its extension to two-dimensional lattices [17, 18]. Motivated by the HN model, we propose a realization of non-Hermiticity in two dimensions exploring a certain resemblance with the Hermitian flux $t$-$\varphi$ model, where $\varphi$ is a staggered magnetic flux [19]. Typically flux phases are constructed by considering hopping matrix elements having the form $t_{\bm{r}_{1},\bm{r}_{2}}=|t_{\bm{r}_{1},\bm{r}_{2}}|e^{i\phi_{\bm{r}_{1},\bm{r}_{2}}}$. For a Hermitian system we require $\phi_{\bm{r}_{1},\bm{r}_{2}}=-\phi_{\bm{r}_{2},\bm{r}_{1}}$. If the latter condition does not hold, we achieve a non-reciprocal hopping akin to the HN model. Our construction starts with a one-dimensional HN model with four sites and periodic boundary conditions having a maximally asymmetric hopping, i.e., one occurring only in one direction, see Fig. 1. Such a simple one-dimensional model is used as a building block for a two-dimensionional non-Hermitian system where each plaquette represents the maximally asymmetric HN model just described. The procedure is schematically illustrated in Fig. 1. By combining these plaquettes in a clock-anticlockwise fashion we arrive at the non-Hermitian version of the flux-like model. Note that the Hermitian $t$-$\varphi$ flux model includes gauge-invariant couplings to the magnetic field, and this can lead to an array of currents giving rise to an antiferromagnetic (AF) flux lattice [19], whereas our model is characterized by asymmetrical hopping and does not include gauge-invariant couplings to an actual magnetic field. It is also interesting to observe that the current pattern following from Fig. 1 bears a resemblance to ice models, more specifically, in the form of a two-dimensional vertex model, which from the figure can be identified to the so called model F [20]. Also worth mentioning in this context is the vertex model associated to the XXZ Heisenberg chain [21], which also takes a non-Hermitian form.
@@ II Model
The proposed generalization of the HN model in two dimensions in the form of the flux model has the following Hamiltonian
$H$ $=$ $-t\sum_{\langle\bm{r},\bm{r}^{\prime}\rangle,\sigma}a^{\dagger}_{\bm{r},\sigma}b_{\bm{r}^{\prime},\sigma}-\mu\sum_{\bm{r},\sigma}f^{\dagger}_{\bm{r},\sigma}f_{\bm{r},\sigma}$ (1)
@@
where
@@
The model is characterized by three parameters, that are the hopping integral $t$, chemical potential $\mu$, and the on-site Hubbard repulsion $U$. The operators $a^{\dagger}_{\bm{r},\sigma}$ and $b_{\bm{r},\sigma}$ are the creation and annihilation operators for fermions with spin $\sigma=\uparrow,\downarrow$ and site index $\bm{r}$ on the sub-lattices $A$ and $B$, respectively, and the hopping processes are shown in Fig. 1. Although the Hamiltonian Eq. (1) is non-Hermitian, it is $\mathcal{PT}$-symmetric. On a lattice, $\mathcal{P}$ and $\mathcal{T}$ operate as follows,
@@
It is instructive first to look into the properties of the non-interacting model where $U=0$. The easiest way to obtain the energy spectrum is to consider the two types of vertices occurring in the lattice of Fig. 1. The bipartite lattice is generated by the blue and red vertices shown in Fig. 2. Thus, picturially, the $\mathcal{PT}$-symmetry is realized by exchanging two vertices ($\mathcal{P}$) and inverting the direction of the arrows ($\mathcal{T}$). The two type of vertices are represented by the operators,
$V_{\bm{r},\sigma}^{A}=a_{\bm{r},\sigma}^{\dagger}(b_{\bm{r}+\hat{\bm{y}},\sigma}+b_{\bm{r}-\hat{\bm{y}},\sigma})+(b_{\bm{r}+\hat{\bm{x}},\sigma}^{\dagger}+b_{\bm{r}-\hat{\bm{x}},\sigma}^{\dagger})a_{\bm{r},\sigma},$ (4)
$V_{\bm{r},\sigma}^{B}=b_{\bm{r},\sigma}^{\dagger}(a_{\bm{r}+\hat{\bm{x}},\sigma}+a_{\bm{r}-\hat{\bm{x}},\sigma})+(a_{\bm{r}+\hat{\bm{y}},\sigma}^{\dagger}+a_{\bm{r}-\hat{\bm{y}},\sigma}^{\dagger})b_{\bm{r},\sigma}.$ (5)
Thus, it is easy to see that,
$H_{U=0}$ $=$ $-\frac{t}{2}\sum_{\bm{r},\sigma}\left(V_{\bm{r},\sigma}^{A}+V_{\bm{r},\sigma}^{B}\right)$ (6)
-- $\mu\left(\sum_{\bm{r}\in A,\sigma}a_{\bm{r},\sigma}^{\dagger}a_{\bm{r},\sigma}+\sum_{\bm{r}\in B,\sigma}b_{\bm{r},\sigma}^{\dagger}b_{\bm{r},\sigma}\right),$
and therefore the energy dispersion is given by $\mathcal{E}^{\pm}_{\bm{k}}=-\mu\pm\varepsilon_{\bm{k}}$ where
$\varepsilon_{\bm{k}}=t\sqrt{\cos k_{x}\cos k_{y}}.$ (7)
For $\mu=0$ (half-filling) the energy spectrum splits into real and imaginary parts, separated by exceptional lines determined by the condition $\cos k_{x}\cos k_{y}=0$. In particular, inside the sublattice Brillouin zone (sBZ), the spectrum is real in the distinct intervals $k_{x}\in(-\pi/2,\pi/2)$ and $k_{y}\in(-\pi/2,\pi/2)$ and there is no gap between the real and imaginary parts of the energies. Remarkably, the exceptional lines coincide here with the Fermi surface as shown in panels (a) and (b) of Fig. 3. For finite $\mu\leq t$ only one of the bands crosses the Fermi level, which is determined by $\mu=\mbox{Re}~{}\varepsilon_{\bm{k}}$ and it does not coincide with exceptional lines. For $\mu>t$ the system is a band insulator.
Figure 3: Calculated energy dispersion of the non-interacting two-dimensional HN flux model for $\mu=0$ (a), $\mu=0.7t$ (c) in the first BZ. Panel (b) shows the Fermi surface $\mu=0$, which are also exceptional lines $\cos k_{x}\cos k_{y}=0$.
@@
$H_{\text{MF}}$ $=$ $-t\sum_{\langle\bm{r},\bm{r}^{\prime}\rangle,\sigma}a^{\dagger}_{\bm{r},\sigma}b_{\bm{r}^{\prime},\sigma}-\mu\sum_{\bm{r},\sigma}f^{\dagger}_{\bm{r},\sigma}f_{\bm{r},\sigma}+\frac{U}{2}\sum_{\bm{r}}m^{2}_{\bm{r}}$ (8)
@@
For $W<t$ the system is an AF metal with complex and real-valued energies depending on the values of the momentum, see Fig. 4-(a,b). The Fermi surface again coincides with exceptional lines that separate real and complex valued energies. For $W\geq t$, the system undergoes a transition from an AF semimetal at $W=t$ (with Dirac points located at (±$(\pm\pi,0)$ and $(0,\pm\pi)$) to an insulator for $W>t$ with real-valued energies Fig. 4-(c), where the $\mathcal{PT}$ symmetry is restored in the entire BZ.
@@
$E_{\bm{k}^{\prime}}^{\pm}=-\mu\pm t\sqrt{\cos^{2}k^{\prime}_{x}+\cos^{2}k^{\prime}_{y}+\frac{W^{2}}{t^{2}}-1},$ (11)
so for $W=t$ the above expression coincides with the spectrum of the $\pi$-flux phase for the square lattice [22, 23]. Moreover, for $W=t$ the energy spectrum $E_{\bm{k}}^{\pm}$ near the four points (±$(\pm\pi,0)$, $(0,\pm\pi)$ in the corners of the sBZ zone is linear $E_{\bm{p}}^{\pm}=-\mu\pm v_{F}|\bm{p}|$, where vector $\bm{p}=(p_{x},p_{y})$ is counted from these four points, and the Fermi velocity is equal to $v_{F}=t/\sqrt{2}$.
Although the spectrum is gapless for the case $W=t$, there is another way to generate a gap. Similarly to graphene [24, 25], we can assume that the two-dimensional system is actually embedded in a three-dimensional space and take into account the long-range Coulomb interaction with a potential $V(\bm{r}-\bm{r}^{\prime})=e^{2}/(\varepsilon|\bm{r}-\bm{r}^{\prime}|)$. In this case, a gap is dynamically generated. This follows by truncating the Schwinger-Dyson equation with the Coulomb interaction as the vertex, which leads to the self-consistent equation,
$\Sigma(\bm{q})=\frac{\alpha}{4\pi}\int d^{2}k\frac{\Sigma(\bm{k})}{\sqrt{\bm{k}^{2}+\Sigma^{2}(\bm{k})}|\bm{k}+\bm{q}|},$ (12)
where $\alpha=e^{2}/(\varepsilon v_{F})$ is the ”fine-structure constant” of the system and $\Sigma(\bm{k})$ is the self-energy. After some simplifications, we can find the following approximate solution to Eq. (12)
$\Sigma(0)=c\Lambda\exp[-\frac{\text{const}}{\sqrt{2\alpha-1}}],$ (13)
where $c$ is a constant and $\Lambda$ is an ultraviolet cutoff. From this we can see that the constant $\alpha$ has a critical value $\alpha_{c}=1/2$ above which the energy spectrum acquires a gap.
Figure 4: Calculated energy spectrum for $\mu=0$ of the AF metal at $W=0.5t$ (a), AF semi-metal at $W=t$ (c) and AF insulator at $W=1.5t$ (d). Panel (b) shows the Fermi surface (exceptional lines) for $W=0.5t$.

### v2

@@ I Introduction
In recent years one of the most popular avenues to study the evolution of open quantum-mechanical systems became its description via effective non-Hermitian Hamiltonians, which break unitarity [1]. At the origin of these research activities lies the seminal work on parity-time ($\mathcal{PT}$) symmetric Hamiltonian with balanced dissipation (loss) and gain energy terms added to Hermitian parent Hamiltonians [2, 3]. $\mathcal{PT}$ symmetry allows to construct further non-Hermitian Hamiltonians, which for some parameter space yield real eigenvalues and the transition between purely real and complex-valued (dissipative) energy spectrum is separated by exceptional points [4, 5]. The field of non-Hermitian systems has grown in recent years, driven by theoretical predictions and experimental discoveries across various field of physics such as optics, acoustics, ultra-cold atomic gases, and superconducting qubits [1, 6, 7]. In particular, non-Hermitian Hamiltonians have recently become quite promising due to the possibility of different realizations, for example using ultracold atoms [8, 9], photonics [10], metamaterials [11, 12] and topological electrical circuits [13, 14, 15]. The characteristic physical phenomena of non-Hermitian systems, such as formation of exceptional points, chiral transport around them, and the non-Hermitian skin effect are typically understood within a single-particle Hamiltonian formalism, and currently the interest shifts towards many-body systems where the interplay between interactions and non-Hermiticity plays a central role.
@@
One prominent example of non-Hermiticity appears in the celebrated Hatano-Nelson (HN) model [16, 17, 18], which consists of a one-dimensional (1D) lattice with asymmetric (non-reciprocal) hoppings, responsible for the localization of all the bulk states at the edges, which leads to the emergence of the so-called topological non-Hermitian skin effect. The non-reciprocity of the hopping may be considered as resulting from some transverse pseudomagnetic field in a cylinder geometry [16]. By now, there has been studies of this model in the presence of interactions [19], disorder [16, 20, 21, 22, 23], as well as its extension to two-dimensional lattices [24, 25]. Motivated by the HN model, we propose a realization of non-Hermiticity in two dimensions exploring a certain resemblance with the Hermitian flux $t$-$\varphi$ model, where $\varphi$ is a staggered magnetic flux [26]. Typically flux phases are constructed by considering hopping matrix elements having the form $t_{\bm{r}_{1},\bm{r}_{2}}=|t_{\bm{r}_{1},\bm{r}_{2}}|e^{i\varphi_{\bm{r}_{1},\bm{r}_{2}}}$. For a Hermitian system we require $\varphi_{\bm{r}_{1},\bm{r}_{2}}=-\varphi_{\bm{r}_{2},\bm{r}_{1}}$. If the latter condition does not hold, we achieve a non-reciprocal hopping akin to the HN model. Our construction starts with a one-dimensional HN model with four sites and periodic boundary conditions having a maximally asymmetric hopping, i.e., one occurring only in one direction, see Fig. 1. Such a simple one-dimensional model is used as a building block for a two-dimensionional non-Hermitian system where each plaquette represents the maximally asymmetric HN model just described. The procedure is schematically illustrated in Fig. 1. By combining these plaquettes in a clock-anticlockwise fashion we arrive at the non-Hermitian version of the flux-like model. Note that the Hermitian $t$-$\varphi$ flux model includes gauge-invariant couplings to the magnetic field, and this can lead to an array of currents giving rise to an antiferromagnetic (AF) flux lattice [26], whereas our model is characterized by asymmetrical hopping and does not include gauge-invariant couplings to an actual magnetic field. It is also interesting to observe that the current pattern following from Fig. 1 bears a resemblance to ice models, more specifically, in the form of a two-dimensional vertex model, which from the figure can be identified to the so called model F [27]. Also worth mentioning in this context is the vertex model associated to the XXZ Heisenberg chain [28], which also takes a non-Hermitian form.
@@ II Model
II.1 Definition of model
The proposed generalization of the HN model in two dimensions in the form of the flux model has the following Hamiltonian,
$H$ $=$ $H_{t}-\mu\sum_{\bm{r},\sigma}f^{\dagger}_{\bm{r},\sigma}f_{\bm{r},\sigma}$ (1)
@@ II.1 Definition of model
where,
$H_{t}=-\frac{1}{4}\sum_{\bm{r},\sigma}\left(V_{\bm{r},\sigma}^{A}+V_{\bm{r},\sigma}^{B}\right),$ (2)
with
$V_{\bm{r},\sigma}^{A}=t_{4}a_{\bm{r},\sigma}^{\dagger}(b_{\bm{r}+\hat{\bm{y}},\sigma}+b_{\bm{r}-\hat{\bm{y}},\sigma})+t_{3}(b_{\bm{r}+\hat{\bm{x}},\sigma}^{\dagger}+b_{\bm{r}-\hat{\bm{x}},\sigma}^{\dagger})a_{\bm{r},\sigma},$ (3)
$V_{\bm{r},\sigma}^{B}=t_{1}b_{\bm{r},\sigma}^{\dagger}(a_{\bm{r}+\hat{\bm{x}},\sigma}+a_{\bm{r}-\hat{\bm{x}},\sigma})+t_{2}(a_{\bm{r}+\hat{\bm{y}},\sigma}^{\dagger}+a_{\bm{r}-\hat{\bm{y}},\sigma}^{\dagger})b_{\bm{r},\sigma}.$ (4)
The model is characterized by three parameters: the hopping integral $t$, chemical potential $\mu$, and the on-site Hubbard repulsion $U$. The operators $a^{\dagger}_{\bm{r},\sigma}$ and $b_{\bm{r},\sigma}$ are the creation and annihilation operators for fermions with spin $\sigma=\uparrow,\downarrow$ and site index $\bm{r}$ on the sub-lattices $A$ and $B$, respectively,
@@
Although the Hamiltonian of Eq. (1) is non-Hermitian, it is $\mathcal{PT}$-symmetric. On a lattice, $\mathcal{P}$ and $\mathcal{T}$ operate as follows,
@@
where $\bm{N}=(N,N)$, with $N$ being the number of lattice sites and $\bm{n}=(1,1)$.
The hopping processes are shown in Fig. 1. In this case we can consider the two types of vertices occurring in the lattice of Fig. 1. The bipartite lattice is generated by the blue and red vertices shown in Fig. 2. Thus, picturially, the $\mathcal{PT}$-symmetry is realized by exchanging two vertices ($\mathcal{P}$) and inverting the direction of the arrows ($\mathcal{T}$). The two type of vertices are represented by the operators, $V_{\bm{r},\sigma}^{A}$ and $V_{\bm{r},\sigma}^{B}$.
II.2 Model in the presence of a magnetic field
Although this model does not include an external magnetic field, we can easily consider its effect by including the gauge-invariant couplings to an actual magnetic field using a Peierls substitution. In this case, the model should be characterized not only by a flux in terms of hopping, but also by a flux in terms of the external magnetic field. We can always choose a gauge in which the phase is equally distributed on all bonds so that $t_{1}=t_{2}=t_{3}^{*}=t_{4}^{*}=te^{-i\varphi}$, where $\varphi=\sum_{\langle\bm{r},\bm{r}^{\prime}\rangle\in p}\varphi_{\bm{r},\bm{r}^{\prime}}$ is the staggered magnetic flux with sum over all plaquettes $p$ and $\varphi_{\bm{r},\bm{r}^{\prime}}=e/\hbar c\int_{\bm{r}}^{\bm{r}^{\prime}}\bm{A}\cdot d\bm{l}$ is the phase, and $\bm{A}$ is the vector potential. It is easy to see that, the energy dispersion for $H_{t}$ is then given by $\mathcal{E}^{\pm}_{\bm{k}}=-\mu\pm\varepsilon_{\bm{k}}$, where
@@ II.2 Model in the presence of a magnetic field
[…]
II.3 Properties of the spectrum
For $U=0$ and $\mu=0$ (half-filling) the energy spectrum splits into real and imaginary parts, separated by exceptional lines determined by the condition $\cos k_{x}\cos k_{y}=0$. In particular, inside the sublattice Brillouin zone (sBZ), the spectrum is real in the distinct intervals $k_{x}\in(-\pi/2,\pi/2)$ and $k_{y}\in(-\pi/2,\pi/2)$ and there is no gap between the real and imaginary parts of the energies. Remarkably, the exceptional lines coincide here with the Fermi surface as shown in panels (a) and (b) of Fig. 3. For finite $\mu\leq t$ only one of the bands crosses the Fermi level, which is determined by $\mu=\mbox{Re}~\varepsilon_{\bm{k}}$ and it does not coincide with exceptional lines. For $\mu>t$ the system is a band insulator.
Figure 3: Calculated energy dispersion of the non-interacting ($W=0$) two-dimensional HN flux model for $\mu=0$ (a), $\mu=0.7t$ (c) in the first BZ. Panel (b) shows the Fermi surface $\mu=0$, which are also exceptional lines $\cos k_{x}\cos k_{y}=0$. Panel (d) shows the Fermi surface for $\mu=0.7$ and exceptional lines.
Figure 4: Calculated energy spectrum of the interacting two-dimensional HN flux model for $\mu=0$ of the AF metal at $W=0.5t$ (a), AF semimetal at $W=t$ (c) and AF insulator at $W=1.5t$ (d). Panel (b) shows the Fermi surface and exceptional lines for $W=0.5t$.
@@ II.3 Properties of the spectrum
$H_{\text{MF}}$ $=$ $H_{t}-\mu\sum_{\bm{r},\sigma}f^{\dagger}_{\bm{r},\sigma}f_{\bm{r},\sigma}+\frac{U}{2}\sum_{\bm{r}}m^{2}_{\bm{r}}$ (8)
@@
For $W<t$ the system is an AF metal with complex and real-valued energies depending on the values of the momentum, see Fig. 4-(a,b). For $\mu=0$ the Fermi surface coincides with exceptional lines that separate real ($\mathcal{PT}$ symmetry is unbroken) and complex ($\mathcal{PT}$ symmetry is broken) valued energies. For $W\geq t$, the system undergoes a transition from an AF semimetal at $W=t$ (with Dirac points located at $(\pm\pi,0)$ and $(0,\pm\pi)$) to an AF insulator for $W>t$ with only real-valued energies Fig. 4-(c,d), where the $\mathcal{PT}$ symmetry is restored in the entire BZ.
@@
$E_{\bm{k}^{\prime}}^{\pm}=-\mu\pm t\sqrt{\cos^{2}k^{\prime}_{x}+\cos^{2}k^{\prime}_{y}+\Delta},$ (11)
where $\Delta=W^{2}/t^{2}-1$ is the gap order parameter. A schematic phase diagram for the metal-insulator phase transition is shown in Fig. 5. For the metallic phase $\Delta<0$ ($W<t$) and the $\mathcal{PT}$ symmetry can be broken for some values of momenta $(k^{\prime}_{x},k^{\prime}_{y})$. For the insulating phase $\Delta>0$ ($W>t$), with the $\mathcal{PT}$ symmetry remaining unbroken for all values of the momenta $(k^{\prime}_{x},k^{\prime}_{y})$. Finally, at the critical value $\Delta=0$ ($W=t$), we have a semimetal phase. The point for which $\Delta=0$ ($W=t$) Eq. (11) coincides with the spectrum of the $\pi$-flux phase for the quantum antiferromagnet in a square lattice [29, 30]. For $W=t$ the energy spectrum $E_{\bm{k}}^{\pm}$ near the four points $(\pm\pi,0)$, $(0,\pm\pi)$ in the corners of the sBZ zone is linear $E_{\bm{p}}^{\pm}=-\mu\pm v_{F}|\bm{p}|$, where vector $\bm{p}=(p_{x},p_{y})$ is counted from these four points, and the Fermi velocity is equal to $v_{F}=t/\sqrt{2}$.
Figure 5: Schematic phase diagram in terms of the gap order parameter $\Delta=W^{2}/t^{2}-1$ distinguishing the AF metal ($\Delta<0$), AF semimetal ($\Delta=0$) and AF insulator ($\Delta>0$).
II.4 Topological triviality of the mean-field Hamiltonian
In this section we briefly discuss the topological properties of our model. To do this we use the matrix representation of the mean-field Hamiltonian in Fourier space
[…]
For such a matrix we can introduce a set of right and left eigenvectors
[…]
where $E_{\bm{k}}^{\pm}$ is given by the Eq. (10). Expressions for the left and right eigenvectors are presented in Appendix B.
Typically, non-trivial topological behavior is associated with a non-zero value of the Chern number. For the Hermitian model, we can define Chern number as [31, 32]
[…]
Using the expressions for left and right eigenvectors that we calculated in Appendix B, we can write
$\hskip-14.22636pt\mathcal{C}^{\pm}$ $=\pm$ $\frac{iWt^{2}}{8\pi}\int_{\text{sBZ}}dk_{x}dk_{y}\frac{\sin k_{x}\sin k_{y}}{(\varepsilon_{\bm{k}}^{2}+W^{2})^{3/2}}=0.$ (15)
The integral is zero because the integrand is odd function with respect to $k_{x}$ and $k_{y}$, and the sBZ is symmetric. This indicates trivial topological properties. However, we should not forget that in non-Hermitian models, phase transitions can occur without changing the Chern number. For example, a phase transition can occur when we pass through an exceptional point, or an exceptional line as in our case.
In addition to the Chern number, we can define for non-Hermitian model the winding number [33, 32, 34]
$\nu^{\pm}=\oint\limits_{S}\frac{d\bm{k}}{2\pi i}\cdot\nabla_{\bm{k}}\log\det[H_{\text{MF}}(\bm{k})-E_{\bm{k}_{\text{EP}}}^{\pm}],$ (16)
where $S$ is a loop in momentum space that encircles one of the exceptional points ($\varepsilon^{2}_{\bm{k}_{\text{EP}}}=-W^{2}$). In our case, $\nu^{\pm}=0$, which again indicates the absence of topological phases.